# Rankify: A Comprehensive Python Toolkit for Retrieval, Re-Ranking, and Retrieval-Augmented Generation

### v2

@@ 4.2. Retrieval Results
Table 4. Performance of re-ranking methods on BM25-retrieved documents for Natural Questions (NQ-Test) and WebQuestions (WebQ-Test). Results are reported in terms of Top-1, Top-5, Top-10, Top-20, and Top-50 accuracy, highlighting the impact of various re-ranking models on retrieval effectiveness.
@@ 4.3. Re-Ranking Results
The performance of various re-ranking methods applied to BM25-retrieved documents is presented in Table 4. The table reports Top-1, Top-5, Top-10, Top-20, and Top-50 accuracy for two datasets: Natural Questions (NQ-Test) and WebQuestions (WebQ). The baseline BM25 retrieval scores serve as a reference, demonstrating the extent to which re-ranking improves document ranking quality across different ranking models.
Across all models, re-ranking consistently improves retrieval effectiveness, with notable gains observed in Top-1 accuracy. Transformer-based ranking models such as UPR (T5-based models), RankGPT, FlashRank (MiniLM and TinyBERT), and RankT5 achieve substantial improvements over the BM25 baseline. In particular, RankT5-3B outperforms other ranking models, achieving a Top-1 accuracy of 47.17% on NQ-Test and 40.40% on WebQ, highlighting the effectiveness of T5-based ranking models for large-scale retrieval. RankGPT, leveraging LLaMA-3.1-8B, also demonstrates strong performance, particularly in WebQ, where it improves Top-1 accuracy from 18.80% (BM25) to 38.77%.
Among lighter-weight models, FlashRank (MiniLM and TinyBERT) and MonoBERT-large provide solid improvements while maintaining efficiency. MiniLM-L-12-v2 achieves 41.02% Top-1 accuracy on NQ-Test and 37.05% on WebQ, making it a strong candidate for deployment scenarios where computational efficiency is critical. The MonoBERT-large model reaches 39.05% on NQ-Test and 34.99% on WebQ, reinforcing the effectiveness of BERT-based cross-encoders for document ranking.
Models that integrate listwise and contrastive ranking techniques, such as Twolar, and LiT5 re-rankers, show competitive performance. Twolar-xl achieves 46.84% Top-1 accuracy on NQ-Test and 41.68% on WebQ, while LiT5-Distill-xl reaches 47.81% on NQ-Test and 42.37% on WebQ, demonstrating that fine-tuned T5 models can significantly improve ranking quality. The sentence-transformer rerankers, particularly GTR-xxl, achieve 42.93% and 39.41% Top-1 accuracy on NQ-Test and WebQ, respectively, making them effective in ranking diverse retrieval outputs.
Overall, the results highlight the importance of re-ranking in improving retrieval effectiveness. While BM25 provides a strong initial retrieval, the application of neural re-rankers significantly enhances retrieval quality, particularly for Top-1 and Top-5 results. The effectiveness of different models varies based on dataset characteristics, with larger transformer-based models excelling in ranking accuracy, while lighter-weight models offer a balance between accuracy and efficiency.
@@ 5. Conclusion
Designed for extensibility, Rankify allows for easy integration of new models and evaluation methods. Future work will focus on improving retrieval efficiency, optimizing re-ranking strategies, and advancing RAG capabilities. The framework is open-source, actively maintained, and freely available for the research community.

### v3

@@ 4.1. Experiment Setup
Table 6. Performance of re-ranking methods on BM25-retrieved documents for NQ Test and WebQ Test. Results are reported in terms of Top-1, Top-5, Top-10, Top-20, and Top-50 accuracy, highlighting the impact of various re-ranking models on retrieval effectiveness. Please note that some results may differ from the original papers (e.g., UPR) as our experiments were conducted with the top 100 retrieved documents, whereas the original studies used 1,000 documents for ranking.
Reranking/ Model NQ WebQ
Top-1 Top-10 Top-50 Top-1 Top-10 Top-50
BM25 - 23.46 56.32 74.57 19.54 53.44 72.34
UPR (Sachan et al., 2022) T5-small 23.60 59.97 75.15 18.55 55.56 72.68
T5-base 26.81 63.32 76.12 20.66 58.56 72.68
T5-large 29.75 65.67 76.48 24.21 60.38 73.32
T0-3B 35.42 67.56 76.75 32.48 64.17 73.67
gpt2 25.95 60.47 75.87 20.47 56.49 72.78
gpt2-medium 26.75 63.04 75.95 22.39 59.54 72.44
gpt2-large 26.59 62.68 75.95 24.06 59.84 72.78
gpt2-xl 27.28 63.24 75.84 23.57 60.48 72.73
gpt-neo-2.7B 28.75 64.81 76.56 24.75 59.64 72.63
RankGPT (Sun et al., 2023) llamav3.1-8b 41.55 66.17 75.42 38.77 62.69 73.12
FlashRank (Damodaran, 2024) TinyBERT-L-2-v2 31.49 61.57 74.95 28.54 60.62 73.17
MultiBERT-L-12 11.99 43.54 69.63 12.54 45.91 67.91
ce-esci-MiniLM-L12-v2 34.70 64.81 76.17 31.84 62.54 73.47
T5-flan 7.95 36.14 66.67 12.05 42.96 67.27
RankT5 (Zhuang et al., 2023) base 43.04 68.47 76.28 36.95 64.27 74.45
large 45.54 70.02 76.81 38.77 66.48 74.31
3b 47.17 70.85 76.89 40.40 66.58 74.45
Inranker (Laitz et al., 2024) small 15.90 46.84 69.83 14.46 46.25 69.98
base 15.90 48.11 69.66 14.46 46.80 69.68
3b 15.90 48.06 69.00 14.46 46.11 69.34
LLM2Vec (BehnamGhader et al., 2024) Meta-Llama-31-8B 24.32 59.55 75.26 26.72 60.48 73.47
MonoBert (Nogueira et al., 2019) large 39.05 67.89 76.56 34.99 64.56 73.96
Twolar (Baldelli et al., 2024) twolar-xl 46.84 70.22 76.86 41.68 67.07 74.40
Echorank (Rashid et al., 2024) flan-t5-large 36.73 59.11 62.38 31.74 58.75 61.51
flan-t5-xl 41.68 59.05 62.38 36.22 57.18 61.51
Incontext Reranker (Chen et al., 2024) llamav3.1-8b 15.15 57.11 76.48 18.89 52.16 71.70
Lit5 (Tamber et al., 2023) LiT5-Distill-base 40.05 65.95 75.73 36.76 63.48 73.12
LiT5-Distill-large 44.40 67.59 76.01 39.66 64.56 73.67
LiT5-Distill-xl 47.81 68.55 76.26 42.37 65.55 73.62
LiT5-Distill-base-v2 42.57 66.73 75.56 39.61 64.22 73.32
LiT5-Distill-large-v2 46.53 67.83 75.87 41.97 65.64 72.98
LiT5-Distill-xl-v2 47.92 69.03 76.17 41.53 65.69 73.27
Sentence Transformer Reranker GTR-base (Ni et al., 2021) 39.41 65.95 76.03 36.56 64.32 73.62
GTR-large 40.63 68.25 76.73 38.97 65.30 73.57
T5-base (Raffel et al., 2020) 31.19 63.60 76.06 29.77 62.84 73.52
T5-large 30.80 63.35 76.37 30.51 61.71 73.37
all-MiniLM-L6-v2 (Wang et al., 2020) 33.35 65.37 76.01 30.95 62.10 73.52
GTR-xl 41.55 67.78 76.81 38.92 66.04 74.01
GTR-xxl 42.93 68.55 77.00 39.41 65.89 74.01
T5-xxl 38.89 67.78 76.64 35.82 65.20 74.01
Bert-co-condensor 30.96 61.91 75.20 32.43 62.20 73.08
Roberta-base-v2 32.60 63.24 75.42 31.34 62.64 73.37
4.2. Retrieval Results
The retrieval performance of BM25 across multiple datasets is shown in Table 3. The results report Top-1, Top-5, Top-10, Top-20, Top-50, and Top-100 retrieval accuracy for each dataset split. Performance varies significantly based on dataset characteristics, with high Top-k accuracy achieved on open-domain QA datasets such as TriviaQA and Natural Questions. On the other hand, more complex multi-hop datasets like 2WikiMultiHopQA and HotpotQA are characterized by lower retrieval rates, reflecting the challenges in retrieving supporting evidence across multiple documents. Temporal datasets like ArchivalQA and ChroniclingAmericaQA exhibit moderate retrieval performance, likely due to the temporal nature of their content. It is important to note that some datasets in Table 3 were either not officially recorded in previous studies or rely on different underlying corpora, making direct comparison difficult. Additionally, some datasets (e.g., ChroniclingAmericaQA) do not report retrieval accuracy as a standard Top-k metric, further complicating direct validation. Despite these variations, our results shown in Table 3 provide a comprehensive overview of BM25’s strengths and limitations.
Table 4 presents the retrieval performance of various dense retrievers across three benchmark datasets: NQ, WebQ, and TriviaQA. The results report Top-1, Top-5, Top-10, Top-20, Top-50, and Top-100 retrieval accuracy, providing insights into how different retrieval models rank relevant documents. Among the tested retrievers, dense retrieval methods such as DPR, MSS-DPR, and ColBERT consistently outperform retrievers like MSS and Contriever. MSS-DPR achieves the highest Top-1 accuracy, reaching 50.16% on NQ, 44.24% on WebQ, and 61.63% on TriviaQA, demonstrating the effectiveness of fine-tuned dense retrieval. Standard DPR also performs well, particularly on NQ and TriviaQA, where it reaches 48.67% and 57.47% Top-1 accuracy, respectively. ColBERT, which employs late-interaction ranking, shows competitive performance, particularly in higher recall settings (Top-50 and Top-100).
Figure 4. Exact Match (EM) for BM25, Contriever, and DPR retrievers across three datasets (NQ, TriviaQA, WebQ) using various language models (LLaMA V3/V3.1, Gemma 2B/9B, LLaMA 2 13B, and Mistral 7B).
4.3. Comparison with Original Implementations
To validate the correctness and reliability of the Rankify’s retrieval module, we compare its performance against original implementations and Pyserini-based baselines for multiple retrievers. Specifically, we evaluate DPR, Contriever, BM25, and ColBERT across three open-domain QA datasets: NQ, TriviaQA, and WebQ. The retrieval performance is measured using Top-20 and Top-100 accuracy.
Table 5 presents the results of our Rankify implementation alongside the original models and Pyserini-based implementations. Our results show that Rankify achieves retrieval performance identical to Pyserini for DPR, BM25 as we utilize Pyserini as the backend for indexing and retrieval. However, for Contriever BGE, and ColBERT, we implemented their methods independently, and our results closely match the original implementations, demonstrating the correctness of our approach.
Additionally, we observe improvements in the performance of BM25 and DPR compared to their original baselines. These improvements can be attributed to the optimizations present in Pyserini, such as better indexing techniques and parameter tuning. For ColBERT and Contriever, we replicate the results reported in their respective papers by running their official GitHub implementations on our datasets, yielding identical outcomes.
@@ 4.4. Re-Ranking Results
The performance of various re-ranking methods applied to BM25-retrieved documents is presented in Table 6. The table reports Top-1, Top-10, and Top-50 accuracy for the NQ-Test and WebQ datasets. The baseline BM25 retrieval scores serve as a reference, demonstrating the extent to which re-ranking improves document ranking quality across different models. Across all methods, re-ranking consistently improves retrieval effectiveness, with notable gains observed in Top-1 accuracy. Transformer-based models such as UPR (T5-based), RankGPT, FlashRank (MiniLM and TinyBERT), and RankT5 achieve substantial improvements over the BM25 baseline. In particular, RankT5-3B outperforms other models, achieving a Top-1 accuracy of 47.17% on NQ-Test and 40.40% on WebQ, highlighting the effectiveness of T5-based models for large-scale retrieval. RankGPT, leveraging LLaMA-3.1-8B, also shows strong performance, particularly in WebQ, where it improves Top-1 accuracy from 18.80% (BM25) to 38.77%.
Among lighter-weight models, FlashRank (MiniLM and TinyBERT) and MonoBERT-large provide solid improvements while maintaining efficiency. MiniLM-L-12-v2 achieves 41.02% Top-1 accuracy on NQ-Test and 37.05% on WebQ, making it a strong candidate for deployment scenarios where computational efficiency is critical. The MonoBERT-large model reaches 39.05% on NQ-Test and 34.99% on WebQ, reinforcing the effectiveness of BERT-based cross-encoders for document ranking. Models that integrate listwise and contrastive ranking techniques, such as Twolar-xl and LiT5, show competitive performance. Twolar-xl achieves 46.84% Top-1 accuracy on NQ-Test and 41.68% on WebQ, while LiT5-Distill-xl reaches 47.81% on NQ-Test and 42.37% on WebQ, demonstrating that fine-tuned T5 models can significantly improve ranking quality. The sentence-transformer rerankers, particularly GTR-xxl, achieve 42.93% and 39.41% Top-1 accuracy on NQ-Test and WebQ, respectively, making them effective for diverse retrieval tasks.
4.5. Generator Results
In this section, we present the performance of Rankify’s generator module when integrated with Retrieval-Augmented Language Modeling (RALM) (Ram et al., 2023). We evaluate the effectiveness of various retrieval models, including BM25, Contriever, and DPR, across three open-domain QA datasets: NQ, TriviaQA, and WebQ. The experiments utilize multiple LLMs: LLaMA V3 8B, LLaMA V3.1 8B, Gemma 2B/9B, LLaMA 2 13B, and Mistral 7B to assess the robustness of the retrieval and generation pipeline. We focus on the Exact Match (EM) metric, which measures the proportion of generated answers that exactly match the ground-truth references. Figure 4 provides a consolidated view of the EM scores for all datasets and models. From the figure, we observe the following key patterns: For the NQ dataset, DPR consistently outperforms other retrievers, achieving its peak performance with the LLaMA V3 8B model (28.08%). On the TriviaQA dataset, BM25 achieves the highest EM score (57.55%) when paired with the Gemma 2 9B model, outperforming both DPR and Contriever. For the WebQuestions (WebQ) dataset, DPR again exhibits strong performance, achieving its best EM score (19.83%) with the LLaMA V3 8B model.
@@ 5. Conclusion
Designed for extensibility, Rankify allows for easy integration of new models and evaluation methods. Future work will focus on improving retrieval efficiency, optimizing re-ranking strategies, and advancing RAG capabilities. The framework is open-source, actively maintained, and freely available for the research community. Future work will focus on incorporating alternative retrieval evaluation metrics such as MAP, Precision, and NDCG, while expanding dataset coverage to include BEIR for broader benchmarking. The framework is open-source, actively maintained, and freely available for the research community.